OpenStudy (anonymous):

Given the function f(x) = x^3 + x^2 − 2x + 1, what is the resulting function when f(x) is shifted to the left 1 unit?

OpenStudy (anonymous):

@kirbykirby @satellite73 I honestly don't know what to do

OpenStudy (anonymous):

Transforming functions is easy f(x) = x^3 + x^2 − 2x + 1 x = af( b(x + c) ) + d . Each letter does something different. c shifts things left or right so the new function, you'll be putting a 1 in the transform it f(x) = (x+1)^3 + (x+1)^2 − (2x+1) + 1

OpenStudy (anonymous):

i have no idea why those weird symbols showed up but ignore them

OpenStudy (anonymous):

\[ f(x + 1) = x^3 + 4x^2 + 3x + 1\]

OpenStudy (anonymous):

f(x) - 1 = x^3 + x^2 - 2x f(x) + 1 = x^3 + x^2 - 2x + 2

OpenStudy (anonymous):

@ohmai which one would it be?

OpenStudy (anonymous):

Its the first one